OpenStudy (anonymous):

Solve for n: 2(3n + 10) = 5(4 - 2n) (4 points) 0 16 all real solutions no solution

OpenStudy (onepieceftw):

Distribute what's in the parentheses 6n+20=20-10n add 10n to both sides 16n+20=20 Subtract 20 from both sides 16n=0 divide both sides by 16 dividing anything by 0 equals 0

OpenStudy (onepieceftw):

Correction diving 0 by anything equals 0

OpenStudy (onepieceftw):

dividing*